# Many-Body Localization in the Age of Classical Computing

### v2

@@ Abstract
This review focuses on recent numerical investigations aiming to clarify the status of the MBL phase, and it establishes the critical open questions about the dynamics of disordered many-body systems. The last decades of research have brought an unprecedented new variety of tools and indicators to study the breakdown of ergodicity, ranging from spectral and wave function measures, observable matrix elements and unitary quantum dynamics, to transport and quantum information measures. We give a comprehensive overview of these approaches and attempt for providing a unified understanding of their main features. We emphasize general trends towards ergodicity with increasing length and time scales, which exclude naive single-parameter scaling hypothesis, necessitate the use of more refined scaling procedures, and which prevent unambiguous extrapolations of numerical results to the asymptotic limit. Providing a concise description of numerical methods for studying ETH and MBL, we explore various approaches to tackle the question of the MBL phase. Persistent finite size drifts towards ergodicity consistently emerge in quantities derived from eigenvalues and eigenvectors of disordered many-body systems. The drifts are related to continuous inching towards ergodicity and non-vanishing transport observed in the dynamics of many-body systems, even at strong disorder. These phenomena impede understanding of microscopic processes at the ETH-MBL crossover. Nevertheless, the abrupt slowdown of dynamics with increasing disorder strength provides premises suggesting the proximity of the MBL phase. This review concludes that the questions about thermalization and its failure in disordered many-body systems remain a captivating area open for further explorations.
@@ I Introduction
Since the XIX century and thanks to the works of Ludwig Boltzmann, James Clerk Maxwell, and Josiah Willard Gibbs, physicists have understood some of the reasons behind the universal and successful applicability of thermodynamics. It was recognized that most of the physical systems undergo a thermalization process in their dynamics, and that their long time behaviour is very accurately described by probability distributions on the phase space, termed statistical ensembles [475, 282]. Time averages of physical observables can then be calculated as ensemble averages with respect to the corresponding statistical ensembles: the statement is commonly referred to as the ergodic hypothesis. The ergodic hypothesis, technically, states that the average over time of all local observables is the same for (almost all) initial states, and therefore it can only depend on the conserved quantities of the system, which for a generic non-integrable system are just macroscopic observables like energy, angular momentum, total magnetization, total charge, etc.
@@
It should be stressed that, while it is widely believed that the ergodic hypothesis holds for the overwhelming majority of many-body systems, the examples of systems in which this statement can be proven are rather rare. On the other hand, there are several examples of systems in which ergodicity breaks down. For instance, systems that undergo second order phase transitions approach, at the critical point, the equilibrium state very slowly, with equilibration times that grow algebraically in the number of microscopic constituents of the system. Typically, however, we term such situations as critical slowing down rather than ergodicity breaking. The situation is much different for disordered systems, such as spin glasses, introduced by [186, 585]. Below the freezing temperature, systems in spin glass phase are trapped in a non-ergodic set of states: the system may fluctuate between several states, but cannot transition to all states of equivalent energy. This was understood within the theory of replica symmetry breaking [468, 469], instrumental for the 2021 Nobel Prize in Physics to G. Parisi [470]. While this solution is exact for the Sherrington-Kirkpatrick model [636], the ergodicity of the Edwards-Anderson model on a finite-dimensional lattice is a more complex problem [623].
@@
Diffusion, which arises in classical systems due to numerous collisions between the constituent particles [189, 615], is one of the standard mechanisms of thermalization in such systems. However, anomalous diffusion is also abundant in classical systems [686]. The recent progress in single particle tracking in biological settings [406] provides one example of anomalous diffusion that can be characterized down to the level of trajectories of individual classical particles [434]. The situation is fundamentally different in the realm of quantum mechanics: diffusion of particles may be inhibited due to presence of disorder, leading to the phenomenon of Anderson localization [27], recognized by the Nobel Prize in Physics for P. Anderson in 1977, and recently observed also in systems of ultracold atoms [75, 524].
@@
Thermalization of quantum systems is typically discussed in terms of quench experiments in which systems are prepared in certain out-of-equilibrium states and then evolve in time. Such a scenario was considered in work [522], which demonstrated that a non-integrable isolated quantum many-body system relaxes to an equilibrium state described by a microcanonical ensemble determined, in turn, by a few variables related to global conserved quantities mentioned before. This process of thermalization must be visible at the level of individual many-body eigenstates, according to the eigenstate thermalization hypothesis (ETH) proposed in the seminal papers [168, 618]. In this sense, an Anderson insulator, as a single particle system, is manifestly non-ergodic since the particles remain close to their original positions rather than propagating throughout the system. While the conditions for the occurrence of Anderson localization in systems of non-interacting quantum particles are well understood, it is of fundamental importance to understand the effects of disorder in a more generic situation – in the presence of interactions between particles.
@@
It was proposed [51, 244, 451] that a sufficiently strong disorder, even in presence of interactions, may inhibit thermalization of quantum many-body systems leading to the phenomenon of many-body localization (MBL). Rather than relaxing to the equilibrium state determined by global conserved quantities, MBL systems retain memory of certain details of the initial state which is preserved in local observables, the so-called local integrals of motion LIOMs (see [293] and references therein) for an indefinite period of time. Such a behavior would result in a transition between an ergodic and MBL phases observable even at infinite temperature, as conjectured in [460].
@@
The existence of MBL transition is of fundamental importance for our understanding of non-equilibrium phenomena in many-body systems, the robustness of diffusive transport, and the applicability of statistical physics. The aim of the present review is to discuss the current status of the phenomenon of MBL and of the putative ETH-MBL phase transition. A particular emphasis will be put on the numerical results which are of paramount importance for our understanding of MBL. The interpretation of the numerical results, as we shall argue, remains to this day inconclusive, despite the intensive efforts of the community interested in non-equilibrium phenomena in quantum many-body systems spanning over the last 15 years. We will show that the difficulties in the interpretation of the numerical data are tied to the exponential growth of many-body Hilbert space with system size $L$, which severely limits system sizes and time scales accessible to present day classical computers and supercomputers. Similar limitations emerge in state-of-the-art experiments, in which perfect isolation from environment can only be maintained for a finite time interval.
@@
The present review is organized as follows. We begin with a concise historical overview of the development of the MBL field in Sec. II. Next, in Sec. III, we provide a definition of MBL and discuss its key phenomenological features. In Sec. IV we describe the current status of the field, highlighting open questions and qualitatively characterizing the challenges associated with interpreting numerical data in strongly disordered interacting systems. We hope that this Section may serve not only as a roadmap of challenges in the field of MBL, but also can be used as a reference point for understanding the difficulties in the interpretation of numerical results for generic complex systems. Sec. V offers a brief exposition of numerical methods typically employed in studies of MBL. In Sec. VI, we provide a quantitative discussion of the ETH-MBL crossover found in the exact diagonalization studies of disordered spin chains, and emphasize the system size drifts, disallowing an unambiguous interpretation of the results. Sec. VII discusses the present understanding of time evolution of disordered many-body systems, highlighting the slow but persistent trends towards thermalization found even at strong disorder. We briefly discuss the relation of thermalization and transport in many-body systems in Sec. VIII, and, in Sec. IX, we outline the microscopic mechanisms responsible for the emergence of ergodicity with decrease of the disorder strength, in particular, the quantum avalanche mechanism and many-body resonances. Finally, in Sec. X we consider various extensions of the MBL phenomenology, while in Sec. XI, we discuss experiments on MBL, highlighting the remaining challenges.
@@ II MBL: historical perspective in short
A seemingly unrelated field of research was conceived in the context of statistical physics and thermalization. In 1955, E. Wigner proposed that statistics of spacings between the lines of heavy atoms nuclei may be modelled by random matrices [708] of appropriate symmetry class, as introduced later by [181]. The observations for the heavy atomic nuclei, which are examples of strongly interacting many-body systems, were followed, among others, by the Bohigas–Giannoni–Schmit conjecture, which asserts that the spectral statistics of quantum systems whose classical counterparts exhibit chaotic behaviour are described by random matrix theory [107, 80], starting the field of quantum chaos [258]. Repulsion of energy levels, characteristic for random matrices, was predicted for electrons in random impurity potentials [21], providing a link between quantum chaos and the physics of disordered quantum systems.
@@
Those results were followed by studies of many-body systems with tools that rely on the random matrix theory [424, 488, 59, 223, 549] and even focused on entanglement in disordered many-body systems [550, 125]. However, the link between quantum chaotic behavior and thermalization of isolated many-body systems, understood in terms of eigenstate thermalization hypothesis (ETH) developed by [168, 618], was established only later, in work [522]. The latter will be further discussed in Sec. III.1.
@@
The developments in the initially unrelated fields of Anderson localization, quantum chaos and thermalization of many-body systems set the stage for studies of many-body localization (MBL), which focus on properties of disordered interacting many-body quantum systems at high temperature. The studies of MBL aim at understanding the impact of interactions and disorder on the dynamics of many body systems, and they typically employ tools of random matrix theory and measures of entanglement from quantum information theory [25, 214] to characterize eigenvalues and eigenvectors of many-body systems. Research on MBL, commenced with the numerical work [451], gained significant momentum when it was postulated that the interplay of disorder and interactions leads to a transition between ergodic and MBL phases that manifests itself in singular changes of many-body eigenstates at arbitrary energies [460], even at infinite temperature. Interest in MBL is additionally driven by the remarkable recent progress in experiments on synthetic quantum matter which allow to directly observe whether and how the many-body systems approach thermal equilibrium [561, 614]. But, despite the experimental progress, numerical simulations on classical computers, that are the focal point of this review, constitute the main source of our understanding of the phenomenon of MBL.
@@ III.1 Quantum thermalization
Statistical mechanics aims to describe complex many-body systems in terms of a few macroscopic quantities. By way of the ergodic hypothesis, one obtains a drastic simplification, going from deterministic equations describing the position and momenta of the classical particles to the statistical properties of such observables. An analogous role for interacting quantum many-body systems is played by the ETH, where the deterministic Schrödinger equation evolution is traded for the statistical properties of matrix elements of local operators.
We consider an isolated quantum many-body system with a Hamiltonian $\hat{H}$ with a complete set of eigenstates $|m\rangle$ corresponding to non-degenerate eigenvalues $E_{m}$, a certain initial state $|\psi\rangle$ that is not an eigenstate of $\hat{H}$, and a local observable $A$. The Schrödinger equation implies that the average value of $A$ after time $\tau$ reads
@@
where $A_{mn}=\langle m|A|n\rangle$ and $c_{n}=\braket{n}{\psi}$. The time average of $A(\tau)$ over time $t$ is given by
@@
Hence, the question about the thermalization of the system, i.e., about the long time behavior of the local observables, is related to properties of the diagonal matrix elements $A_{mm}$. Moreover, the second term in (1) illustrates the time scale of the approach to the equilibrium value, $\lim_{t\to\infty}\bar{A}(t)$, which is determined by the properties of the off-diagonal matrix elements $A_{mn}$. This time scale can be, in some cases, exponentially long in the system size as it is determined by the differences of energies in the middle of many-body spectrum.
Early works [168, 618] paved the way to formulation of ETH as an ansatz for both diagonal and off-diagonal matrix elements of observables in the eigenbasis of $\hat{H}$ [619] [^1] See Sec. III of [489] and Sec. 4.2 of [148] for survey of works reporting results consistent with Eq. (3) obtained in the semi-classical limit of quantum systems whose classical counterpart is chaotic.,
@@
where $\bar{E}=(E_{m}+E_{n})/2$ is the mean energy, $\omega_{mn}=E_{m}-E_{n}$ is the energy difference, and $R_{mn}$ is a random variable with zero mean and unit variance. The function $S(\bar{E})$ is the thermodynamic entropy at energy $\bar{E}$, i.e., the logarithm of the density of states [101], and $\mathcal{A}(\bar{E})$ and $f_{\mathcal{A}}(\bar{E},\omega)$ are smooth functions of their arguments. Within the random matrix theory (RMT), $\mathcal{A}(\bar{E})={\rm const}$ and $f_{\mathcal{A}}(\bar{E},\omega)=1$. A nontrivial dependence on $\bar{E}$ in $\mathcal{A}(\bar{E})$ can be attributed to the presence of conserved operators, including the Hamilton operator and its higher powers [420], while a nontrivial dependence on $\omega$ in $f_{\mathcal{A}}(\bar{E},\omega)$ contains information on the nature of transport and the fine structure of response functions [148]. Indeed, in quantum many-body systems, the function $f_{\mathcal{A}}(\omega)$ (we skip the $\bar{E}$ dependence here) typically interpolates between the characteristic plateau $f_{\mathcal{A}}(\omega)=\mathrm{const}$ at small frequency $\omega$ and a large frequency decay $f_{\mathcal{A}}(\omega)\stackrel{{\scriptstyle\omega\to\infty}}{{\longrightarrow}}0$. The change of the behavior of $f_{\mathcal{A}}(\omega)$ between the two limits is defines a Thouless energy $\omega_{\mathrm{Th}}$, and the relaxation processes of physical observables are expected to be completed at the Thouless time $t_{\mathrm{Th}}\propto\omega_{\mathrm{Th}}^{-1}$. The scaling of $t_{\mathrm{Th}}$ with system size $L$ may determine the type of dynamics and transport in the system, for instance, for diffusive dynamics one expects $t_{\mathrm{Th}}\propto L^{2}$. The notions of Thouless time and Thouless energy in disordered many-body systems are discussed more quantitatively in Secs. VI and VII, respectively.
Expanding (3) for the diagonal matrix elements around the mean energy $\braket{E}=\braket{\psi}{\hat{H}}{\psi}$, assuming that the energy variance $\delta E^{2}$ in the initial state |$\ket{\psi}$ is sufficiently small, one arrives at [522]
@@
with corrections that typically decay algebraically with the number of lattice sites. Equation (4) shows that the equilibrium value of a local observable $A$ is an average over a microcanonical ensemble of states $\hat{\rho}_{MC}$, and $\mathcal{N}_{\langle E\rangle,\delta E}$ determines the number of states in this ensemble. This is the notion of quantum ergodicity: all eigenstates |$\ket{m}$ of $\hat{H}$ in the energy shell $|E_{m}-\langle E\rangle|<\delta E$ contribute equally to the long time average value of $A$, in close analogy to the classical statistical ensembles that include all microstates compatible with macroscopic constraints with equal probability.
It is nowadays well established that the ETH represents a sufficient criterion for the onset of thermalization [148]. The systems that satisfy ETH typically exhibit diffusive transport; to which extent the ETH can be extended to systems with non-diffusive types of transport, in particular when the system is open to the boundary, is discussed in Sec. IV.1. Other recent activities in the studies of ETH include, among other things, the role of the matrix elements correlations for higher point correlation functions [465], and its implications for the sensitivity of eigenstates to perturbations [463].

### v3

@@ Abstract
This review focuses on recent numerical investigations aiming to clarify the status of the MBL phase, and it establishes the critical open questions about the dynamics of disordered many-body systems. The last decades of research have brought an unprecedented new variety of tools and indicators to study the breakdown of ergodicity, ranging from spectral and wave function measures, observable matrix elements, and unitary quantum dynamics, to transport and quantum information measures. We give a comprehensive overview of these approaches and attempt to provide a unified understanding of their main features. We emphasize general trends towards ergodicity with increasing length and time scales, which exclude naive single-parameter scaling hypothesis, necessitate the use of more refined scaling procedures, and prevent unambiguous extrapolations of numerical results to the asymptotic limit. Providing a concise description of numerical methods for studying ETH and MBL, we explore various approaches to tackle the question of the MBL phase. Persistent finite size drifts towards ergodicity consistently emerge in quantities derived from eigenvalues and eigenvectors of disordered many-body systems. The drifts are related to continuous inching towards ergodicity and non-vanishing transport observed in the dynamics of many-body systems, even at strong disorder. These phenomena impede the understanding of microscopic processes at the ETH-MBL crossover. Nevertheless, the abrupt slowdown of dynamics with increasing disorder strength provides premises suggesting the proximity of the MBL phase. This review concludes that the questions about thermalization and its failure in disordered many-body systems remain a captivating area open for further explorations.
@@ I Introduction
Since the XIX century and thanks to the works of Ludwig Boltzmann, James Clerk Maxwell, and Josiah Willard Gibbs, physicists have understood some of the reasons behind the universal and successful applicability of thermodynamics. It was recognized that most of the physical systems undergo a thermalization process in their dynamics, and that their long time behaviour is very accurately described by probability distributions on the phase space, termed statistical ensembles [513, 301]. Time averages of physical observables can then be calculated as ensemble averages with respect to the corresponding statistical ensembles: the statement is commonly referred to as the ergodic hypothesis. Technically, the ergodic hypothesis states that the long-time averages of few-body observables are the same for (almost all) initial states. Hence, these averages can only depend on the conserved quantities of the system, which, in a generic non-integrable system, are just macroscopic observables such as energy, angular momentum, total magnetization, total charge, etc.
@@
It should be stressed that, while it is widely believed that the ergodic hypothesis holds for the overwhelming majority of many-body systems, the examples of systems in which this statement can be proven are rather rare. Systems that undergo second order phase transitions approach, at the critical point, the equilibrium state very slowly, with equilibration times that grow algebraically in the number of microscopic constituents of the system. Typically, however, we term such situations as critical slowing down rather than ergodicity breaking. The situation is much different for disordered systems, such as spin glasses, introduced by [196, 632]. Below the freezing temperature, systems in spin glass phase are trapped in a non-ergodic set of states: the system may fluctuate between several states, but cannot transition to all states of equivalent energy. This was understood within the theory of replica symmetry breaking [506, 507], instrumental for the 2021 Nobel Prize in Physics to G. Parisi [508]. While this solution is exact for the Sherrington-Kirkpatrick model [687], the ergodicity of the Edwards-Anderson model on a finite-dimensional lattice is a more complex problem [673].
@@
Diffusion, which arises in classical systems due to numerous collisions between the constituent particles [199, 665], is one of the standard mechanisms of thermalization in such systems. However, a phenomenon that is also abundant in classical systems is anomalous diffusion [741]. The recent progress in single particle tracking in biological settings [442] provides one example of anomalous diffusion that can be characterized down to the level of trajectories of individual classical particles [472]. The situation is fundamentally different in the realm of quantum mechanics: diffusion of particles may be inhibited due to presence of disorder, leading to the phenomenon of Anderson localization [29], recognized by the Nobel Prize in Physics for P. Anderson in 1977, and recently observed also in systems of ultracold atoms [78, 569].
@@
Thermalization of quantum systems is typically discussed in terms of quench experiments in which systems are prepared in certain out-of-equilibrium states and then evolve in time. Such a scenario was considered in [566], which demonstrated that a non-integrable isolated quantum many-body system relaxes to an equilibrium state described by a microcanonical ensemble determined, in turn, by a few variables related to global conserved quantities mentioned before. This process of thermalization must be visible at the level of individual many-body eigenstates, according to the eigenstate thermalization hypothesis (ETH) proposed in the seminal papers [177, 668]. In this sense, an Anderson insulator, as a single particle system, is manifestly non-ergodic since the particles remain close to their original positions rather than propagating throughout the system. While the conditions for the occurrence of Anderson localization in systems of non-interacting quantum particles are well understood, it is of fundamental importance to understand the effects of disorder in a more generic situation – in the presence of interactions between particles.
@@
It was proposed [54, 260, 489] that a sufficiently strong disorder, even in presence of interactions, may inhibit thermalization of quantum many-body systems leading to the phenomenon of many-body localization (MBL). Rather than relaxing to the equilibrium state determined by global conserved quantities, MBL systems retain memory of certain details of the initial state that is preserved in local observables, the so-called local integrals of motion LIOMs, see [318] and references therein, for an indefinite period of time. Such a behavior would result in a transition between an ergodic and MBL phases observable even at infinite temperature, as conjectured in [498].
@@
The existence of MBL transition is of fundamental importance for our understanding of non-equilibrium phenomena in many-body systems, the robustness of diffusive transport, and the applicability of statistical physics. The aim of the present review is to discuss the current status of the phenomenon of MBL and of the putative ETH-MBL phase transition. A particular emphasis will be put on the numerical results which are of paramount importance for our understanding of MBL. The interpretation of many numerical results, as we shall argue, remains to this day inconclusive, despite the intensive efforts of the community interested in non-equilibrium phenomena in quantum many-body systems spanning over the last 15 years. An abrupt slow down of many-body dynamics with increasing disorder strength makes it possible to distinctly observe an MBL regime, in which finite size systems fail to thermalize even over extended periods of time. However, the question of whether the observed MBL regime actually corresponds to an MBL phase – a dynamical phase of matter in which the thermalization does not occur at arbitrarily long times and across all system sizes – remains unresolved. The difficulties in the interpretation of the numerical results and understanding the status of the MBL phase are tied to the exponential growth of many-body Hilbert space with system size $L$, which severely limits system sizes and time scales accessible to present day classical computers and supercomputers. Similar limitations emerge in state-of-the-art experiments, in which perfect isolation from environment can only be maintained for a finite time interval.
@@
The present review is organized as follows. We begin with a concise historical overview of the development of the MBL field in Sec. II. Next, in Sec. III, we provide a definition of MBL and discuss its key phenomenological features. In Sec. IV we describe the current status of the field, highlighting open questions and qualitatively characterizing the challenges associated with interpreting numerical results in strongly disordered interacting systems. We hope that this Section may serve not only as a roadmap of challenges in the field of MBL, but also can be used as a reference point for understanding the difficulties in the interpretation of numerical results for generic complex systems. Sec. V offers a brief exposition of numerical methods typically employed in studies of MBL. In Sec. VI, we provide a quantitative discussion of the ETH-MBL crossover found in the exact diagonalization studies of disordered spin chains, and emphasize the system size drifts, disallowing an unambiguous interpretation of the results. Sec. VII discusses the present understanding of time evolution of disordered many-body systems, highlighting the slow but persistent trends towards thermalization found even at strong disorder. We briefly discuss the relation of thermalization and transport in many-body systems in Sec. VIII, and, in Sec. IX, we outline the microscopic mechanisms responsible for the emergence of ergodicity with decrease of the disorder strength, in particular, the quantum avalanche mechanism and many-body resonances. Finally, in Sec. X we consider various extensions of the MBL phenomenology, while in Sec. XI, we discuss experiments on MBL, highlighting the remaining challenges.
@@ II MBL: historical perspective in short
A seemingly unrelated field of research was conceived in the context of statistical physics and thermalization. In 1955, E. Wigner proposed that statistics of spacings between the lines of heavy atomic nuclei may be modelled by random matrices [764] of appropriate symmetry class, as introduced later by [191]. The observations for the heavy atomic nuclei, which are examples of strongly interacting many-body systems, were followed, among others, by the Berry conjecture, which suggested that quantum states associated with irregular stochastic classical motion appear to be Gaussian random function [65], and by the Bohigas–Giannoni–Schmit conjecture, which asserts that the spectral statistics of quantum systems whose classical counterparts exhibit chaotic behaviour are described by random matrix theory [110, 83]. These early contributions started the field of quantum chaos [274]. Repulsion of energy levels, characteristic for random matrices, was predicted for electrons in random impurity potentials [23], providing a link between quantum chaos and the physics of disordered quantum systems.
@@
Those results were followed by studies of many-body systems with tools that rely on the random matrix theory [460, 527, 62, 239, 595] and even focused on entanglement in disordered many-body systems [596, 162]. However, the link between quantum chaotic behavior and thermalization of isolated many-body systems, understood in terms of eigenstate thermalization hypothesis (ETH) developed by [177, 668], was established in the context of systems that do not have semi-classical limit, e.g., spin-1/2 models, only later, in [566] [^1] See [325] for an early study of applicability of statistical mechanics to dynamics of spin-1/2 chains.. The latter will be further discussed in Sec. III.1.
@@
The developments in the initially unrelated fields of Anderson localization, quantum chaos and thermalization of many-body systems, set the stage for studies of many-body localization (MBL), which focus on properties of disordered interacting many-body quantum systems at high temperature. The studies of MBL aim at understanding the impact of interactions and disorder on the dynamics of many body systems, and they typically employ tools of random matrix theory and measures of entanglement from quantum information theory [27, 228] to characterize eigenvalues and eigenvectors of many-body systems. Research on MBL, commenced with the numerical study [489], gained significant momentum when it was postulated that the interplay of disorder and interactions leads to a transition between ergodic and MBL phases that manifests itself in singular changes of many-body eigenstates at arbitrary energies [498], even at infinite temperature [^2] Throughout this review, the notions of infinite and high temperature are used interchangeably and, unless otherwise noted, mean the energy density corresponding to the infinite temperature.. Interest in MBL is additionally driven by the remarkable recent progress in experiments on synthetic quantum matter which allow to directly observe whether and how the many-body systems approach thermal equilibrium [608, 664]. But, despite the experimental progress, numerical simulations on classical computers, that are the focal point of this review, constitute the main source of our understanding of the phenomenon of MBL.
@@ III.1 Quantum thermalization
Statistical mechanics aims to describe complex many-body systems in terms of a few macroscopic quantities. By way of the ergodic hypothesis and mixing property, one obtains a drastic simplification, going from deterministic equations describing the position and momenta of classical particles to the statistical properties of such observables. An analogous role for interacting quantum many-body systems is played by the ETH, where the deterministic Schrödinger equation evolution is traded for the statistical properties of matrix elements of operators.
We consider an isolated quantum many-body system governed by a Hamiltonian $\hat{H}$ with a complete set of eigenstates $|m\rangle$ corresponding to non-degenerate eigenvalues $E_{m}$, an initial state $|\psi\rangle$ that is not an eigenstate of $\hat{H}$, and an observable $\hat{A}$ which is few-body, i.e., acting on number of particles (or spins) much smaller than the total number of constituents of the system. We set $\hbar=1$. The Schrödinger equation implies that the average value of $\hat{A}$ after time $\tau$ reads
@@
where $A_{mn}=\langle m|\hat{A}|n\rangle$ and $c_{n}=\braket{n}{\psi}$. The time average of $A(\tau)$ over time $t$ is given by
@@
Hence, the question about the the thermalization of the system, i.e., about the long time behavior of the observables, is related to properties of the diagonal matrix elements $A_{mm}$. Moreover, the second term in (1) sets the time scale of the approach to the equilibrium value, $\lim_{t\to\infty}\bar{A}(t)$, which is determined by the properties of the off-diagonal matrix elements $A_{mn}$. This time scale can be, in some cases, exponentially long in the system size as it is determined by the differences of energies in the middle of many-body spectrum333 In many cases, the approach to the equilibrium value is much faster, see, e.g., [562]. .
Early works [177, 668] paved the way to formulation of ETH as an ansatz for both diagonal and off-diagonal matrix elements of observables in the eigenbasis of $\hat{H}$ [669] [^4] See Sec. III of [528] and Sec. 4.2 of [153] for survey of works reporting results consistent with Eq. (3) obtained in the semi-classical limit of quantum systems whose classical counterpart is chaotic, including [520, 213, 214]. See also [693] for Boltzmann distribution derivation using typicality arguments.,
@@
where $\bar{E}=(E_{m}+E_{n})/2$ is the mean energy, $\omega_{mn}=E_{m}-E_{n}$ is the energy difference, and $R_{mn}$ is a random variable with zero mean and unit variance. The function $S(\bar{E})$ is the thermodynamic entropy at energy $\bar{E}$, i.e., the logarithm of the density of states [104], and $\mathcal{A}(\bar{E})$ and $f_{\mathcal{A}}(\bar{E},\omega)$ are smooth functions of their arguments. Within the random matrix theory (RMT), $\mathcal{A}(\bar{E})={\rm const}$ and $f_{\mathcal{A}}(\bar{E},\omega)=1$. A nontrivial dependence on $\bar{E}$ in $\mathcal{A}(\bar{E})$ can be attributed to the presence of conserved operators, including the Hamilton operator and its higher powers [281, 456], while a nontrivial dependence on $\omega$ in $f_{\mathcal{A}}(\bar{E},\omega)$ contains information on the nature of transport and the fine structure of response functions [153]. Indeed, in quantum many-body systems, the function $f_{\mathcal{A}}(\omega)$ (we skip the $\bar{E}$ dependence here) typically interpolates between the characteristic plateau $f_{\mathcal{A}}(\omega)=\mathrm{const}$ at small frequency $\omega$ and a large frequency decay $f_{\mathcal{A}}(\omega)\stackrel{{\scriptstyle\omega\to\infty}}{{\longrightarrow}}0$. The change of the behavior of $f_{\mathcal{A}}(\omega)$ between the two limits defines the Thouless energy $\omega_{\mathrm{Th}}$, and the relaxation processes of physical observables are expected to be completed at the Thouless time $t_{\mathrm{Th}}\propto\omega_{\mathrm{Th}}^{-1}$. The scaling of $t_{\mathrm{Th}}$ with system size $L$ may determine the type of dynamics and transport in the system, for instance, for diffusive dynamics one expects $t_{\mathrm{Th}}\propto L^{2}$. The notions of Thouless time and Thouless energy in disordered many-body systems are discussed more quantitatively in Secs. VI and VII, respectively.
Expanding (3) for the diagonal matrix elements around the mean energy $\braket{E}=\braket{\psi}{\hat{H}}{\psi}$, and assuming that the energy variance $\delta E^{2}$ in the initial state |$\ket{\psi}$ is sufficiently small, one arrives at [566]
@@
with corrections that typically decay algebraically with the number of lattice sites. Equation (4) shows that the equilibrium value of a few-body observable $\hat{A}$ is an average over a microcanonical ensemble of states $\hat{\rho}_{MC}$, and $\mathcal{N}_{\langle E\rangle,\delta E}$ determines the number of states in this ensemble. This is the notion of quantum ergodicity: the long time average of the observable $\hat{A}$ is independent of the specific arrangement of the coefficients $|c_{m}|^{2}$, and is the same as the average of $A_{mm}$ over the entire energy shell $|E_{m}-\langle E\rangle|<\delta E$. This closely parallels the classical statistical ensembles that include all microstates compatible with macroscopic constraints with equal probability.
It is nowadays well established that the ETH represents a sufficient criterion for the onset of thermalization [153]. The systems that satisfy ETH typically exhibit diffusive transport; to which extent the ETH can be extended to systems with non-diffusive types of transport, in particular when the system is open to the boundary, is discussed in Sec. IV.1. Other recent activities in the studies of ETH include, among others, the role of the matrix elements correlations for higher point correlation functions [222, 503], and the modification of the ETH in the vicinity of the ergodicity breaking transition point [356].